OpenStudy (logan13):

Janine inflated 1 ball to a radius of 18 cm and another ball to a radius of 12 cm. How much more air was in the larger ball? Use 3.14 to approximate pi and express your answer in hundredths. cm3

OpenStudy (justdoublechecking):

Volume of a sphere = 4/3*pi*r³

OpenStudy (justdoublechecking):

1.3333333*3.14* 18³= ? 1.3333333*3.14* 12³= ?

OpenStudy (logan13):

hmm

OpenStudy (justdoublechecking):

Subtract answers to find the difference between them.

OpenStudy (justdoublechecking):

Did you get the answer yet?

OpenStudy (logan13):

No it does not show up in the caculator

OpenStudy (justdoublechecking):

I see... I don't know then... It worked for me.

OpenStudy (logan13):

i forgot to include a number

OpenStudy (justdoublechecking):

Did it work this time?

OpenStudy (logan13):

766.159980846

OpenStudy (justdoublechecking):

What was your formula? My answer is way different.

OpenStudy (logan13):

1.3333333*3.14* 18³= thats what i put in

OpenStudy (justdoublechecking):

Hold on...

OpenStudy (justdoublechecking):

I got 24,422.91721

OpenStudy (logan13):

how do i put that in

OpenStudy (logan13):

the answer

OpenStudy (justdoublechecking):

It is that^ minus the answer you got for the other ball.

OpenStudy (justdoublechecking):

(7,236.4199)

OpenStudy (logan13):

how would i put in the second thing

OpenStudy (logan13):

in my caculator

OpenStudy (justdoublechecking):

Answer A-Answer B= Final Answer

OpenStudy (logan13):

@JustDoubleChecking 17186.49731

OpenStudy (justdoublechecking):

Yup. Thats what I got! Is that one of your options?

OpenStudy (logan13):

if thats the answer i have to put in down in a box

OpenStudy (logan13):

thank you

OpenStudy (justdoublechecking):

Oh. Ok. Yes. I believe that is your answer.

OpenStudy (justdoublechecking):

Your welcome. Good job!